How many kilos containing 6% butterfat must be added to 220 kilos of cream containing 22% butterfat to make a mixture containing 10% butterfat?
Answer by stanbon(75887)

How many kilos containing 6% butterfat must be added to 220 kilos of cream containing 22% butterfat to make a mixture containing 10% butterfat?
--------------
Equation:
fat + fat = fat
-----
0.06x + 0.22*220 = 0.10(x+220)
------------
6x + 22*220 = 10x + 10*220
-4x = -12*220
x = 3*220
x = 660 kilos (amt of 6% needed)
Cheers,
Stan H.
----------------